OpenStudy (anonymous):

Add the proper constant to each binomial so that the resulting trinomial is a perfect square trinomial. Then factor the trinomial: x^2+x+____

OpenStudy (anonymous):

Look at your coefficient for the x term. Call that b. You always want your constant term to be (1/2 * b)^2. In your example b is 1

OpenStudy (anonymous):

think of it this way (x+k)^2 = x^2 + 2kx + k^2 we want 2k = 1 thus k = 1/2 so (1/2)^2 = 1/4 so you want to add 1/4 x^2 + x + 1/4 = (x + 1/2)^2